OpenStudy (anonymous):

radical 75x^4 y^5? simplify

OpenStudy (anonymous):

lets go @saifoo.khan ??

OpenStudy (anonymous):

i dont think those are radicals

OpenStudy (saifoo.khan):

Like that?

OpenStudy (anonymous):

yeess

OpenStudy (anonymous):

\[\sqrt{75x^4 y^5} \implies \sqrt{3 \times 5 \times 5 \times x^2 \times x^2 \times y^2 \times y^2 \times y} = 5x^2 y^2\sqrt{3y}\]

OpenStudy (anonymous):

is this corect @waterineyes

OpenStudy (anonymous):

@saifoo.khan ????

OpenStudy (saifoo.khan):

@broderick365 : Too many cooks spoil the broth.

OpenStudy (saifoo.khan):

And @broderick365 that's correct. (:

OpenStudy (anonymous):

thansk!!

OpenStudy (anonymous):

Why I am not getting any notification and no mentions ??